OpenStudy (anonymous):

find the derivative of:

OpenStudy (anonymous):

\[\sqrt{x}secx+3\]

OpenStudy (anonymous):

\[\sqrt{x} secx \tan x + \frac{ 1 }{ \sqrt{x} }secx\]

OpenStudy (anonymous):

product rule \[(fg)'=f'g+g'f\]with \[f(x)=\sqrt{x},f'(x)=\frac{1}{2\sqrt{x}}, g(x)=\sec(x),g'(x)=\sec(x)\tan(x)\]

OpenStudy (anonymous):

missing a 2 in the denominator of the second term

OpenStudy (anonymous):

\[\sqrt{x}secxtanx+\frac{ 1 }{ 2\sqrt{x} } secx\] Thanks satellite73!